OpenStudy (anonymous):

A 20 kg cart is moving at 2.0 m/s when a 5.0 N force is applied for a distance of 6.0 m in the direction the car is moving. What is the kinetic energy of the cart at the end of the 6.0 m?

OpenStudy (anonymous):

@undeadknight26 ? Any chance you could help me out? :)

undeadknight26 (undeadknight26):

o.o I learned this stuff a while ago so i don;t really member it...

OpenStudy (anonymous):

Oh that's too bad :/ Any idea who might know? Someone else I could ask...?

OpenStudy (anonymous):

@Loser66 ? Is there any chance you know how to do this...?

OpenStudy (anonymous):

Or @wolfe8 ?

OpenStudy (anonymous):

I'm assuming it's just 1/2 * 20 * 2.0^2

OpenStudy (loser66):

you have to find out V by \( V^2 = V_0^2 + 2as\) to get Velocity. to find a , you need apply F = ma where F = 5.0N , m = 20kg after then, put all into kinetic energy formula K.E = 1/2 mv ^2

OpenStudy (anonymous):

So a = 25 What is s? seconds?

OpenStudy (anonymous):

*0.25

OpenStudy (loser66):

s is the length it moves = 6.0 m

OpenStudy (anonymous):

2.0^2 + 3 = V^2 ?

OpenStudy (anonymous):

V^2 = 7?

OpenStudy (loser66):

yup

OpenStudy (anonymous):

So 70 J would be my answer

OpenStudy (anonymous):

Great! Could you help me with one more? (I'll start a new thread)